### v2

# Covariate-Adaptive Randomization in Clinical Trials without Inflated Variances

Abstract
Covariate adaptive randomization (CAR) procedures are extensively used to reduce the likelihood of covariate imbalances occurring in clinical trials. In literatures, a lot of CAR procedures have been proposed so that the specified covariates are balanced well between treatments. However, the variance of the imbalance of the unspecified covariates may be inflated comparing to the one under the simple randomization. The variance inflation causes the usual test of treatment effects being not valid and adjusting the test being not an easy work. In this paper, we propose a new kind of covariate adaptive randomization procedures to balance covariates between two treatments with a ratio $\rho:(1-\rho)$. Under this kind of CAR procedures, the convergence rate of the imbalance of the specified covariates is $o(n^{1/2})$, and at the same time the asymptotic variance of the imbalance of any unspecified (observed or unobserved) covariates does not exceed the one under the simple randomization. The ”shift problem” found by Liu, Hu, and Ma (2025) will not appear under the new CAR procedures.
Keywords: Balancing covariates; Covariate adaptive randomization; Simple randomization; Pocock and Simon’s procedure; Clinical trial; shift problem
AMS 2010 subject classifications: Primary 60F05; Secondary 62G10
1 Introduction
Covariate adaptive randomization (CAR) procedures have often been implemented for balancing treatment allocation over influential covariates in clinical trials (McEntegart, , 2003; Rosenberger and Lachin, , 2015; Taves, , 2010; Lin, Zhu and Su, , 2015). The main purposes are to enhance the comparability of treatment groups and to increase the efficiency of the statistical inference. Many CAR procedures have been proposed to balance treatment assignments within strata and over covariate margins, provided that the covariates under consideration are discrete (categorical with two or more levels) (Pocock and Simon, , 1975; Antognini and Zagoraiou, , 2011; Hu and Hu, , 2012; Rosenberger and Lachin, , 2015). The theories on CAR randomization procedures are established progressively in the past decade(Hu and Hu, , 2012; Ma, Hu and Zhang, , 2015; Hu and Zhang, , 2020; Hu, Ye and Zhang, , 2023). For balancing continuous covariates, besides the method of discretization, a variety of methods have been proposed by minimization of prescribed imbalance measures (Hoehler, , 1987; Nishi and Takaichi, , 2004; Endo et al., , 2006; Stigsby and Taves, , 2010; Ciolino et al., , 2011; Lin and Su, , 2012; Ma and Hu, , 2013; Qin et al., , 2018; Jiang, Ma and Yin, , 2018; Li, Zhou and Hu, , 2019). Recently, Ma et al, (2024) unified a lot of CAR procedures by proposing a general CAR procedure for balancing specified covariate features $\bm{\phi}(\bm{X}_{i})$ between two treatments $1$ and $2$, where $\bm{\phi}(\bm{X}_{i})=(\phi_{1}(\bm{X}_{i}),\ldots,\phi_{q}(\bm{X}_{i}))$ can contain additional covariate features besides the covariate vector $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ of a unit in clinical trals, such as quadratic and interaction terms. It is shown that the convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ is $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$, where $n$ is the sample size, $\bm{X}_{i}$ is the covariate vector of the $i$-th unit, and $T_{i}=1$, 0 stands for the $i$-unit being allocated to treatment $1$ and $2$, respectively. Under some suitable conditions on the density of $\bm{\phi}(\bm{X}_{i})$, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ can attain the fastest convergence rate $O_{P}(1)$, and at the same time, for any unspecified covariate feature $m(\bm{X}_{i})$, the asymptotic normality of its normalized imbalance $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ holds and the asymptotic variance $\sigma_{m}^{2}$ is nonzero when $m(\bm{X}_{i})$ is not a linear function of $\bm{\phi}(\bm{X}_{i})$. However, $\sigma_{m}^{2}$ may exceed $\textsf{E}[m^{2}(\bm{X}_{i})]/4$ that is the variance under the complete randomization in which each unit is allocated to treatments $1$ and $2$ with equal probability $1/2$. This issue is referred to as the variance inflation. The variance inflation under the procedure of Pocock and Simon, (1975) is also founded by Zhao et al., (2024). Because the asymptotic variance of the test of treatment effects is usually a function of $\sigma_{m}^{2}$ for some unspecified covariate feature $m(\bm{X}_{i})$, the variance inflation causes such test being not valid (Ma et al, , 2024; Zhang, , 2023). The closed form of $\sigma_{m}^{2}$ is also unknown so that estimating it and adjusting the test are not easy works. A lot of methods for the adjustment to obtain robust statistical inferences in the CAR clinical trials have being proposed(Bugni, Canay and Shaikh, , 2018; Ye and Shao, , 2020; Liu and Yang, , 2020; Ma, Tu and Liu, , 2024; Wang et al., , 2023). The purpose of this paper is to propose a new kind of CAR procedures so that the specified covariate features $\bm{\phi}(\bm{X}_{i})$ can be balanced well and at the same time the asymptotic variance $\sigma_{m}^{2}$ of the normalized imbalance of any unspecified covariate feature $m(\bm{X}_{i})$ is not inflated and has a closed form.
From the proofs (Hu and Zhang, , 2020; Ma et al, , 2024; Zhang, , 2023), we can find that the convergence of $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ and the finiteness of its asymptotic variance $\sigma_{m}^{2}$ are strongly based on the theory of Markov chains and the property that the allocation probabilities are symmetric about treatments so that, as a Markov chain, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ has a symmetric invariant distribution and thus $(T_{i}-1/2)m(\bm{X}_{i})$s have zero means under the invariance distribution for any $m(\bm{X}_{i})$. Recently, Liu, Hu and Ma, (2025) generalized Ma et al, (2024)’s procedure to balance covariances between two treatments with a ratio $\rho:(1-\rho)$ instead of half to half. The convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ of the specified covariate feature $\bm{\phi}(\bm{X}_{i})$ is also shown to be $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$. However, by their theory, for an unspecified covariate feature $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/n$ may converge to a nonzero constant $c_{m}$ due to the fact that $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is no longer a symmetric Markov chain when $\rho\neq 1/2$. This issue is refer to as ”shift problem” of covariate $m(\bm{X}_{i})$. The shift problem is a bad news because in testing hypotheses in the covaraite-adaptive randomized clinical trials, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=o_{P}(n)$ is a basic condition (Ma, Hu and Zhang, , 2015; Ma et al., , 2020; Ma et al, , 2024; Zhu et al, , 2025; Liu and Hu, , 2023; Liu, Hu and Ma, , 2025), and sometimes the convergence rate $O_{P}(\sqrt{n})$ is needed (Ma, Tu and Liu, , 2024). We will show that our proposed CAR procedures are valid for all the case of $0<\rho<1$. Under the proposed CAR procedure, the ”shift problem” will no longer appear and the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ will be always controlled by $\rho(1-\rho)\textsf{E}[m^{2}(\bm{X}_{i})]$, the variance under the simple randomization in which each unit is randomized to treatments $1$ and $2$ respectively with $\rho$ and $1-\rho$, for any $m(\bm{X}_{i})$.
In the next section, the framework of the new CAR procedures are proposed with theories on its properties. In Section 3, as an application, the properties of a simple test of treatment effects under the proposed CAR procedure are studied based on the observed responses on treatments. It is shown that the test always controls the type I error rate without any information of the covariates, and the test can be adjusted to have the precise the type I error rate and increase the power when the data of the covariate features $\bm{\phi}(\bm{X}_{i})$s in the randomization stage are also available in the analysis stage. The proofs are given in the last section.
2 New family of CAR procedures
2.1 Framework
Suppose that $n$ experimental units are to be assigned to two treatment groups. Let $T_{i}$ be the assignment of the $i$-th unit, i.e., $T_{i}=1$ for treatment 1 and $T_{i}=0$ for treatment 0. Consider $p$ covariates available for each unit and use $\bm{X}_{i}=(X_{i,1},...,X_{i,p})\in\mathbb{R}^{p}$ to denote the covariates of the $i$-th unit. Besides $\bm{X}_{i}$, the unit may have other random features such as responses and other covaraites which are observable or unobservable and not considered in the randomization, we denote them by $Z_{i},W_{i}$. We assume that $\{(\bm{X}_{i},Z_{i},W_{i});i=1,2,\ldots,n\}$ are independent and identically distributed as a random vector $(\bm{X},W,Z)$. We consider a procedure to balance general covariate features $\bm{\phi}(\bm{X}_{i})$, defined by a feature map $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\to\mathbb{R}^{q}$ that maps $\bm{X}_{i}$ into a $q$-dimensional feature space. Here, $q$ is usually larger than $p$ so that $\bm{\phi}(\bm{X}_{i})$ has more features than the original covariates. The vector $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is the imbalance measure of the covariate feature $\bm{\phi}(\bm{X})$, where $0<\rho<1$. The numerical imbalance measure $Imb_{n}$ is defined as the squared Euclidean norm of the imbalance vector $\bm{\Lambda}_{n}$, i.e.,
$Imb_{n}=\Big\|\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})\Big\|^{2}.$
In this paper, we denote $\langle\bm{x},\bm{y}\rangle=\bm{x}\bm{y}^{\prime}=\sum_{t=1}^{d}x_{i}y_{j}$, $\|\bm{x}\|=\sqrt{\langle\bm{x},\bm{x}\rangle}$, $\bm{x}^{\otimes 2}=\bm{x}^{\prime}\bm{x}=\big(x_{i}x_{j}\big)_{d\times d}$, for real vectors $\bm{x}=(x_{1},\ldots,x_{d})$ and $\bm{y}=(y_{1},\ldots,y_{d})$.
Notice that
$Imb_{n}-Imb_{n-1}=$ $\|\bm{\Lambda}_{n}\|^{2}-\|\bm{\Lambda}_{n-1}\|^{2}$
$=$ $2(T_{n}-\rho)\langle\bm{\Lambda},\phi(\bm{X}_{n-1})\rangle+(T_{n}-\rho)^{2}\|\bm{\phi}(\bm{X}_{n})\|^{2}.$
When the $n$-th unit is ready to be randomized to treatments, the values of $\bm{\Lambda}_{n-1}$ and $\bm{X}_{n}$ are available and given. The decreasing or increasing of the numerical imbalance measure $Imb_{n-1}$ depends on the value of $\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle$ and the allocation probability. We define the allocation probability as a decreasing function of $\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle$ so that when the value of $|\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle|$ is large, the value of $Imb_{n}$ will be drifted back. The new $\bm{\phi}$-based CAR procedure is defined as follows:
  1. (1)
    Randomly assign the first unit to treatment $1$ with the probability $\rho$ and to treatment $2$ with the probability $1-\rho$.
  2. (2)
    Suppose that (n1)(n-1) units have been assigned to a treatment $(n>1)$ and the $n$-th unit is to be assigned, and the results of assignments $T_{1},\ldots,T_{n-1}$ of previous stages and all covariates $\bm{X}_{1},\ldots,\bm{X}_{n}$ up to the $n$-th unit are observed. Calculate the imbalance vector $\bm{\Lambda}_{n-1}$.
  3. (3)
    Assign the $n$-th unit to the treatment with the probability
    $\ell_{n}=:\textsf{P}(T_{n}=1|\bm{X}_{n},...,\bm{X}_{1},T_{n-1},...,T_{1})=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle}{(n-1)^{\gamma}}\right),$ (2.1)
    where $0<\gamma<1$, $\ell(x):\mathbb{R}\to(0,1)$ is a non-increasing function with
    $\ell(0)$ $=\rho,\;\ell^{\prime}(0)<0\text{ and }\ell(x)\text{ is twice differentiable at }x=0.$ (2.2)
  4. (4)
    Repeat the last two steps until all units are assigned.
The function $\ell(x)$ in (2.1) is called the allocation function. A proposed choice of $\ell$ is
$\ell(x)=\frac{\big[2\rho\Phi(-x)\big]\wedge 1+1-\big[2(1-\rho)\Phi(x)\big]\wedge 1}{2}.$ (2.3)
where $\Phi$ is the standard normal distribution, $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$. Another simple choice of $\ell$ is
$\ell(x)=\Phi(-x+u_{\rho}),$ (2.4)
[…]
According to the theoretical results given in Theorem 2.2 and Remark 2.2, the parameter $\gamma$ controls the size of the imbalance measure $Imb_{n}$. The smaller the parameter $\gamma$ , the smaller is $Imb_{n}$, but possibly the larger is the imbalance $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}$ of the unspecified feature $Z_{i}$ that is not a linear fuction of $\bm{\phi}(\bm{X}_{i})$. The value of $\gamma$ is suggested to be in $[0.5,1)$ so that the asymptotic expectation of $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}/n$ does not exceed the one under the simple randomization for any $Z_{i}$s.
2.2 Theoretical Properties
In this subsection we shall study the theoretical properties of the CAR procedures with a feature map $\bm{\phi}(\bm{x})=(\phi_{1}(\bm{x}),\ldots,\phi_{q}(\bm{x})):\mathbb{R}^{p}\mapsto\mathbb{R}^{q}$. Our goal is to obtain the convergence rate of covariate imbalance measured by $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(2T_{i}-1)\bm{\phi}(\bm{X}_{i})$ and the asymptotic normality of $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}$.
To obtain the asymptotic properties, we require the following assumptions.
Assumption 2.1.
The sequence $\{(\bm{X}_{i},Z_{i},W_{i})\}$ is a sequence of independent and identically distributed as a random vector $(\bm{X},W,Z)$..
Assumption 2.2.
The feature map $\phi(\bm{X})$ satisfies that $\textsf{E}[\|\phi(\bm{X})\|^{r_{0}}]$ is finite, $r_{0}>2$.
In the sequel of this paper, for sequences of real numbers $\{a_{n}\}$, $\{b_{n}\}$ $(b_{n}>0)$ and a sequence of random vectors $\{\bm{Y}_{n}\}$, $a_{n}=O(b_{n})$ means $|a_{n}/b_{n}|\leq C$ for some $C>0$, $a_{n}=o(b_{n})$ means $a_{n}/b_{n}\to 0$, $\bm{Y}_{n}=O(b_{n})$ (resp. $o(b_{n})$) in $L_{r}$ means $(\textsf{E}\|\bm{Y}_{n}\|^{r})^{1/r}=O(b_{n})$ (resp. $o(b_{n})$), $\bm{Y}_{n}=o_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}\to 0$ in probability, and $\bm{Y}_{n}=O_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}$ is bounded in probability.
Theorem 2.1.
Suppose that Assumptions 2.1 and 2.2 are satisfied. Then
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\gamma}+n^{\frac{1}{2}+\frac{1}{2(r_{0}-1)}}\big)=o(n).$ (2.5)
If $\textsf{E}[|Z|]<\infty$ and $\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}$, then
$\textsf{E}|(\ell_{n}-\rho)Z_{n}|\to 0,$ (2.6)
$\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\right|}{n}\to 0.$ (2.7)
Moreover, if $r_{0}\geq 4$, $\gamma>\frac{1}{2}+\frac{1}{4(r_{0}-1)}$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}[W]=0$, then
$\left(\frac{\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}}{\sqrt{n}},\frac{\sum_{i=1}^{n}W_{i}}{\sqrt{n}}\right)\overset{d}{\to}N(0,\vec{\sigma}_{Z}^{2})\times N(0,\textsf{E}W^{2}),$ (2.8)
where $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[\widetilde{Z}^{2}]$, $\widetilde{Z}=Z-\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ and $\bm{x}_{0}=\arg\min\{\textsf{E}\big[(Z-\langle\bm{\phi}(\bm{X}),\bm{x}\rangle)^{2}\big]:\bm{x}\in\mathbb{R}^{q}\}$.
Remark 2.1.
(i) No shift problem:
Liu, Hu and Ma, (2025) proposes a CAR procedure for balancing covariate features $\bm{\phi}(\bm{X}_{i})$ between treatments with an unequal ratio $\rho:(1-\rho)$. When the additional covariate $m(\bm{X}_{i})$ is not the linear function of $\bm{\phi}(\bm{X}_{i})$, it may happen that $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}c\neq 0$. This issue is referred as ”shift problem”. The simulation results presented in the Supplementary Material of Liu, Hu and Ma, (2025) show that the shift problem on the additional covariates is common when balancing continuous covariates by using their CAR procedure. Now, (2.7) shows that the shift problem no longer exists under our proposed CAR procedure.
(ii) Closed form of variance:
$\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ is the orthogonal projection of $Z$ on $\bm{\phi}(\bm{X})$, we denote it by $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]$. Thus, $\vec{\sigma}_{Z}^{2}=\textsf{E}\big(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}$. This closed form makes $\vec{\sigma}_{Z}^{2}$ can be estimated basing on the observations of $Z$ and $\bm{\phi}(\bm{X})$.
It is easily seen that $\textsf{E}\big(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}=\textsf{E}Z^{2}-\textsf{E}\big(\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}\leq\textsf{E}Z^{2}$. Thus, $\vec{\sigma}_{Z}^{2}=0$ if and only if $Z\in Span\{\bm{\phi}(\bm{X})\}=:\{\langle\bm{x},\bm{\phi}(\bm{X})\rangle:\bm{\in}\mathbb{R}^{q}\}$, and $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[Z^{2}]$ if and only if $\textsf{E}[\bm{\phi}(\bm{X})Z]=\bm{0}$, i.e., $Z\perp\bm{\phi}(\bm{X})$. As a result, $\vec{\sigma}_{Z}^{2}=\textsf{Var}(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})])+\vec{\sigma}_{\textsf{E}[Z|\bm{\phi}(\bm{X})]}^{2}$. Further, if $1\in Span\{\bm{\phi}(\bm{X})\}$, then $\vec{\sigma}_{Z}^{2}=\vec{\sigma}_{Z-\textsf{E}Z}\leq\textsf{Var}(Z)$.
(iii)No variance inflation:
The simple randomization is randomizing each patient to treatment $1$ with the same probability $\rho$. For the simple randomization, $\vec{\sigma}_{Z}^{2}=\rho(1-\rho)\textsf{E}Z^{2}$. So, the proposed CAR procedure always has smaller asymptotic variance than the simple randomization for all observed or unobserved feature $Z_{i}$ with $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\neq 0$. The variance inflation is no longer a problem under our proposed CAR procedure.
(iv) Independence of exogenous variables:
The normalized imbalance $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}/\sqrt{n}$ of $Z_{i}$s is an endogenous variable of the design, and the normalized sum $\sum_{i=1}^{n}W_{i}/\sqrt{n}$ of $W_{i}$s can be regraded as an exogenous variable. (2.8) means that they are asymptotic independent regardless $W_{i}$ and $Z_{i}$ themselves are independent or not.
When $\bm{\phi}(\bm{X})$ has finite moments of all orders. The convergence rate of the imbalance $\bm{\Lambda}_{n}$ can be improved a lot.
Theorem 2.2.
Suppose Assumption 2.1, and that 2.2 is satisfied for all $r_{0}>2$.
(2.6) and(2.7) hold if $\textsf{E}[|Z|]<\infty$, and
$\big(\textsf{E}\|\bm{\Lambda}_{n}\|^{r}\big)^{1/r}=O(n^{\gamma/2})\;\text{ for all }r>0.$ (2.9)
If $\textsf{E}Z^{2}<\infty$, then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\big((T_{i}-1)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]\big)$
$+O(n^{\gamma/2}+n^{1-\gamma})\text{ in }L_{r}$ (2.10)
for all $r>0$, where $\widetilde{Z}_{i}=Z_{i}-\textsf{Pj}[Z_{i}|\bm{\phi}(\bm{X}_{i})]$.
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$ and $1/2<\gamma<1$, then (2.8) holds,
$\frac{\textsf{E}\Big(\sum_{i=1}^{n}(T_{i}-\rho)Z_{i},\sum_{i=1}^{n}W_{i}\Big)^{\otimes 2}}{n}\to diag(\vec{\sigma}_{Z}^{2},\textsf{E}W^{2}),$ (2.11)
and
$\textsf{E}\Big(\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\Big)^{2}\leq C_{\phi}\textsf{E}Z^{2}\text{ with }C_{\phi}\text{ does not depend on }Z.$ (2.12)
Remark 2.2.
If $\ell(x)$ is thrice differentiable at $x=0$, and $\ell^{\prime\prime}(0)=0$ or $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$, then $n^{1-\gamma}$ in ((ii)) can be repalced by $n^{1-3\gamma/2}$, and (2.8), (2.11), (2.12) remain true when $1/3<\gamma<1$.
The allocation function $\ell$ as chosen in (2.4) with $\rho=1/2$ or (2.3) satisfies $\ell^{\prime}(0)=-\frac{1}{\sqrt{2\pi}}$ and $\ell^{\prime\prime}(0)=0$.
If the distribution of $(\bm{\phi}(\bm{X}),Z)$ is symmetric, i.e. $(\bm{\phi}(\bm{X}),Z)$ and $(\bm{\phi}(\bm{X}),Z)$ have the same distribution, then $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$.
Finally, we propose an allocation function such that (2.8) and (2.11) hold for all $0<\gamma<1$. The function is defined by
$\ell(x)=\ell_{propose}(x)=\Phi(-x+u_{\rho/2})\vee(\rho-\lambda x)\wedge\Phi(-x+u_{(\rho+1)/2}),$ (2.13)
where $\lambda>0$ is a tuning parameter. In (2.13), $\Phi(-x+u_{\rho/2})$, $\Phi(-x+u_{(\rho+1)/2})$ can be replaced by $\alpha\rho$, $1-(1-\alpha)\rho$, respectively, $0<\alpha<1$.
Theorem 2.3.
Suppose Assumption 2.1, and that 2.2 is satisfied for all $r_{0}>2$, and $\textsf{E}Z^{2}<\infty$. Let $\ell(x)$ be defined as in (2.13). Then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}\big((T_{i}-1)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]\big)+O(n^{\gamma/2})\text{ in }L_{r}$ (2.14)
for all $r>0$. Moreover, (2.8), (2.11) and (2.12) hold for all the case of $0<\gamma<1$ if $\textsf{E}W^{2}<\infty$ and $\textsf{E}W=0$.
From the proofs given in Section 4, the conclusions in Theorems 2.1-2.3 remain true when the factor $(n-1)^{\gamma}$ in the definition (2.1) of the allocation probability is replaced by a general factor $\lambda_{n-1}$ with $0<c_{0}\leq\lambda_{n}/n^{\gamma}\leq C_{0}<\infty$. Thus, $\bm{\Lambda}_{n-1}$ and $\bm{\phi}(\bm{X}_{n})$ in (2.1) can be normalized by $c_{1}\vee\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\|\bm{\phi}(\bm{X}_{i})\|^{2}}\wedge C_{2}$, respectively, to eliminate scale sensitivity. Here, $c_{1},C_{2}$ are positive constants which are needed in the technical proofs. In practices, we can choose a small $c_{1}$ and a large $C_{2}$.
Example 2.1.
(Discrete covariates). Suppose $\bm{X}=(X_{1},\ldots,X_{p})$ consists of $p$ discrete covariates, with the $l$-th covariate $X_{l}$ having $m_{l}$ levels, $x_{l}^{j},j=1,\cdots,m_{l}$, $l=1,\cdots,p$. We let $\bm{\phi}=\bm{\phi}(\bm{X})$ be the following feature map in our framework,
$\bm{\phi}=\left(\sqrt{w_{o}},\underset{\sum m_{l}\text{marginal terms}}{\underbrace{\ldots,\sqrt{w_{m,l}}I\{X_{l}=x_{l}^{k_{l}^{\ast}}\},\ldots}},\underset{\prod m_{l}\text{within-stratum terms}}{\underbrace{\ldots,\sqrt{w_{s}}I\{X_{1}=x_{1}^{k_{1}^{\ast}},\ldots,X_{p}=x_{p}^{k_{p}^{\ast}}\}\ldots}}\right),$
where $w_{o}$, $w_{m,l}$, $w_{s}$ are the corresponding nonnegative weights and at least one of them is no zero.
Let $D_{n}=\sum_{i=1}^{n}(T_{i}-\rho)$, $D_{n}(l,k_{1}^{\ast})=(T_{i}-\rho)I\{X_{i,l}=x_{l}^{k_{l}^{\ast}}\}$, $D_{n}(k_{1}^{\ast},\ldots,k_{p}^{\ast})=(T_{i}-\rho)I\{X_{i,1}=x_{1}^{k_{1}^{\ast}},\ldots,X_{i,p}=x_{p}^{k_{p}^{\ast}}\}$. Then
$\|\bm{\Lambda}_{n}\|^{2}=w_{0}D_{n}^{2}+\sum_{l=1}^{p}w_{m,l}\sum_{k_{l}=1}^{m_{l}}D_{n}^{2}(l;k_{l})+w_{s}\sum_{k_{1},\ldots,k_{p}}D_{n}^{2}(k_{1},\ldots,k_{p}).$
When $\bm{X}_{n}=(x_{1}^{k_{1}^{\ast}},\ldots,x_{p}^{k_{p}^{\ast}})$,
$\Lambda_{n-1}(k_{1}^{\ast},\ldots,k_{p}^{\ast})=:$ $\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle$
$=$ $w_{0}D_{n-1}+\sum_{l=1}^{p}w_{m,l}D_{n-1}(l;k_{l}^{\ast})+w_{s}D_{n-1}(k_{1}^{\ast},\ldots,k_{p}^{\ast}).$
Then, the allocation probability of the $n$th-unit is
$\textsf{P}\Big(T_{n}=1$ $\Big|\bm{X}_{1},\ldots,\bm{X}_{n-1},T_{1},\ldots,T_{n-1},\bm{X}_{n}=(x_{1}^{k_{1}^{\ast}},\ldots,x_{l}^{k_{p}^{\ast}})\Big)$
$=\ell\left(\frac{\Lambda_{n-1}(k_{1}^{\ast},\ldots,k_{p}^{\ast})}{(n-1)^{\gamma}}\right).$ (2.15)
By Theorem 2.3,
$\|\bm{\Lambda}_{n}\|^{2}=O(n^{\gamma})\text{ in }L_{r}\text{ for all }r>0.$ (2.16)
Thus, when $w_{s}\neq 0$, $D_{n}^{2}(k_{1},\ldots,k_{p})=O_{P}(n^{\gamma/2})$ and so for any $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=O_{P}(n^{\gamma/2})$. When $w_{s}=0$, we will have
$D_{n}(l,k_{l})=$ $O_{P}(n^{\gamma/2}),\text{ if }w_{m,l}\neq 0,$
$D_{n}=$ $O_{P}(n^{\gamma/2}).$
In any case, we have
$\frac{\textsf{E}(\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i}))^{2}}{n}\to\vec{\sigma}_{m}^{2}\text{ and }\frac{\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})}{\sqrt{n}}\overset{d}{\to}N(0,\vec{\sigma}_{m}^{2})$ (2.17)
with $\vec{\sigma}_{m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}+\vec{\sigma}_{\textsf{E}m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}\leq\rho(1-\rho)\textsf{Var}(m(\bm{X}))$ since $1\in Span(\bm{\phi})$. In particular,
$\frac{D_{n}(k_{1},\ldots,k_{p})}{\sqrt{n}}\overset{d}{\to}N(0,\vec{\sigma}^{2}(k_{1},\ldots,k_{p}))$ (2.18)
with $\vec{\sigma}^{2}(k_{1},\ldots,k_{p})\leq\rho(1-\rho)p(k_{1},\ldots,k_{p})(1-p(k_{1},\ldots,k_{p}))$, where $p(k_{1},\ldots,k_{p})=\textsf{P}(X_{1}=x_{1}^{k_{1}},\ldots,X_{p}=x_{p}^{k_{p}})$.
When $\rho=1/2$, the allocation probability is similar to that of Zhang, (2021), where (2.16) is obtained, but the properties (2.17) and (2.18) are not founded.
When $\rho=1/2$ and $\gamma=0$, our CRA procedure is similar to that of Hu and Hu, (2012). The theory of this kind of CAR procedures can be found in Hu and Zhang, (2020) and Zhang, (2023). It is shown that (2.17) still holds, but $\vec{\sigma}_{m}^{2}$ has no closed from and may exceed $\textsf{E}[m^{2}(\bm{X})]/4$ unless $\textsf{E}[m(\bm{X})|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$. In particular, when $w_{s}=w_{0}=0$, the procedure is Pocock and Simon, (1975)’s marginal procedure. Zhao et al., (2024) demonstrated via examples that along some directions, there is an increased variance of within-stratum imbalances under Pocock and Simon, (1975)’s procedure compared to the simple randomization.
Example 2.2.
When a linear model is adopted to fit the responses of treatments over a covariate feature $\bm{X}=(X_{1},\ldots X_{p})$. Adopting the feature $\bm{\phi}=(1,X_{1},\ldots,X_{p})$ in our framework will yields
$\sum_{i=1}^{n}(T_{i}-\rho)=O(n^{\gamma})\text{ and }\sum_{i=1}^{n}(T_{i}-\rho)X_{i,t}=O(n^{\gamma})\text{ in }L_{r},$
when $X_{1},\ldots,X_{p}$ have finite moments of each order. At the same time, for any nonlinear function $m=m(X_{1},\ldots,X_{p})$ with finite variance, then the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ is $\rho(1-\rho)\textsf{Var}(m-\textsf{Pj}[m|\bm{X}])$ which is controlled by $\rho(1-\rho)\textsf{Var}(m)$. The procedure will not cause the ”shift problem” and the variance inflation.
The results in Theorems 2.1-2.3 does not include the case of $\gamma=0$. Though $\gamma=0$ is not our proposed choice, we give the results about this case as follows.
Theorem 2.4.
Let $\gamma=0$. Suppose that Assumptions 2.1 and 2.2 are satisfied. We have the following conclusions.
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\frac{1}{r_{0}-1}}\big)$, and $O\big(n^{\frac{8}{(r_{0}+1)^{2}}}\big)$ when $r_{0}\geq 3$.
If $\textsf{E}|Z|<\infty$ and $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, then (2.7) holds and
$\textsf{E}\left(\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\right)^{2}\leq 2n\textsf{E}Z^{2}+C_{0}\textsf{E}Z^{2}\textsf{E}\|\bm{\Lambda}_{n}\|^{2}\text{ if }\textsf{E}Z^{2}<\infty.$ (2.19)
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$, $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}\in Span\{\bm{\phi}(\bm{X})\}$ when $\rho\neq 1/2$, then
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}}{\sqrt{n}}\overset{d}{\to}N(0,\vec{\sigma}_{Z}^{2}),$ (2.20)
where $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}]$. Further, if $\textsf{E}W^{2}<\infty$, $\textsf{E}[W]=0$ and also $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)W$ is in $Span\{\bm{\phi}(\bm{X})\}$, then we have (2.8).
When $\gamma=0$, the CAR procedure is similar to Liu, Hu and Ma, (2025)’s procedure. It does not cause the “shift problem” when $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and the variance is not inflated when $\rho=1/2$. When $\bm{X}$ is discrete as in Example 2.1 and $\bm{\phi}(\bm{X})$ given there with $w_{s}\neq 0$, i.e., the stratified randomization is adopted, then $\textsf{E}[Z|\bm{\phi}(\bm{X})]$, $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}$, $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)W$ are all in $Span\{\bm{\phi}(\bm{X})\}$, and so the asymptotic normality (2.8) holds with a variance $\vec{\sigma}_{Z}^{2}$ being not inflated. This conclusion is consistent with Theorem 3 of Liu, Hu and Ma, (2025).
3 An application
As an application, we study hypothesis testing of treatment effects under the proposed covariate-adaptive procedure. For each unit $i$, denote $Y_{i}(1)$ and $Y_{i}(2)$ as the potential outcomes under the treatments $1$ and $2$, respectively, with $\textsf{E}[Y_{i}(t)]=\mu_{t}$ and $\textsf{Var}(Y_{i}(t))=\sigma_{t}^{2}$, $t=1,2$. The observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(2).$ The treatment effect is defined as $\tau=\textsf{E}\{Y_{i}(1)\}-\textsf{E}\{Y_{i}(0)\}$. Also, denote $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ as the covariate of unit $i$. Assume that $(Y_{i}(0),Y_{i}(1),\bm{X}_{i})$, $i=1,\ldots,n$, are independent and identically distributed as $(Y(0),Y(1),\bm{X})$. We suppose that the units are randomized sequentially by the proposed CAR procedure with feature $\bm{\phi}(\bm{X})$. We consider the test of
$H_{0}:\tau=:\textsf{E}\{Y_{i}(1)\}-\textsf{E}\{Y_{i}(2)\}=0\;\;\text{versus}\;\;H_{A}:\tau\neq 0.$ (3.21)
Since $\overline{Y}_{n,1}=\frac{1}{N_{n,1}}\sum_{i=1}^{n}T_{i}Y_{i}$, $\overline{Y}_{n,2}=\frac{1}{N_{n,2}}\sum_{i=1}^{n}(1-T_{i})Y_{i}$, $\widehat{\sigma}_{n,1}^{2}=\frac{1}{N_{n,1}-1}\sum_{i=1}^{n}T_{i}(Y_{i}-\overline{Y}_{n,1})^{2}$, $\widehat{\sigma}_{n,2}^{2}=\frac{1}{N_{n,2}-1}\sum_{i=1}^{n}(1-T_{i})(Y_{i}-\overline{Y}_{n,2})$ are consistent estimators of $\mu_{1}$, $\mu_{2}$, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$, respectively, where $N_{n,1}=\sum_{i=1}^{n}T_{i}$ and $N_{n,2}=n-N_{n,1}$, we consider the classical test statistic
$\mathcal{T}(n)=\frac{\overline{Y}_{n,1}-\overline{Y}_{n,2}}{\sqrt{\widehat{\sigma}_{n,1}^{2}/N_{n,1}+\widehat{\sigma}_{n,2}^{2}/N_{n,2}}}.$ (3.22)
Theorem 3.1.
Suppose that the proposed CAR procedure with a feature $\bm{\phi}(\bm{X})$ is applied to balance covariates with Assumptions 2.1, 2.2 for all $r_{0}>2$ being satisfied. Then
$\overline{Y}_{n,t}\overset{P}{\to}\mu_{t},\;\;\widehat{\sigma}_{n,t}^{2}\overset{P}{\to}\sigma_{t}^{2},\;\;t=1,2,$ (3.23)
$\sqrt{n}\big(\overline{Y}_{n,1}-\overline{Y}_{n,2}-\tau\big)\overset{D}{\to}N(0,\sigma_{\tau}^{2}),$ (3.24)
where
$\sigma_{\tau}^{2}=\frac{\sigma_{1}^{2}}{\rho}+\frac{\sigma_{2}^{2}}{1-\rho}-\rho(1-\rho)\sigma_{Pj}^{2}=\frac{\widetilde{\sigma}_{1}^{2}}{\rho}+\frac{\widetilde{\sigma}_{2}^{2}}{1-\rho},$ (3.25)
$\sigma_{Pj}^{2}=$ $\textsf{E}\Big(\textsf{Pj}\big[\frac{Y(1)-\mu_{1}}{\rho}+\frac{Y(2)-\mu_{2}}{1-\rho}\big|\bm{\phi}(\bm{X})\big]\Big)^{2},$
$\widetilde{\sigma}_{t}^{2}=$ $\textsf{E}\Big(Y(t)-\mu_{t}-\textsf{Pj}\big[Y(t)-\mu_{t}\big|\bm{\phi}(\bm{X})\big]\Big)^{2}$
$+\rho(1-\rho)\textsf{E}\Big(\textsf{Pj}\big[Y(1)-\mu_{1}-(Y(2)-\mu_{2})\big|\bm{\phi}(\bm{X})\big]\Big)^{2},\;t=1,2.$
Moreover,
  • (i)
    Under $H_{0}:\tau=0$,
    $\mathcal{T}(n)\overset{D}{\to}N(0,\sigma_{\mathcal{T}}^{2}),\;\text{with},\;\sigma_{\mathcal{T}}^{2}=\frac{\sigma_{\tau}^{2}}{\sigma_{e}^{2}}\leq 1,$ (3.26)
    where $\sigma_{e}^{2}=\frac{\sigma_{1}^{2}}{\rho}+\frac{\sigma_{2}^{2}}{1-\rho}$.
  • (ii)
    under $H_{A}:\tau\neq 0$, consider a sequence of local alternatives, i.e., $\tau=\delta/\sqrt{n}$ for a fixed $\delta\neq 0$ (we suppose the distributions of $Y(t)-\mu_{t}$, $t=1,2$, do not depend on $n$), then
    $\mathcal{T}(n)\overset{D}{\to}N(\Delta,\sigma_{\mathcal{T}}^{2}),\;\text{with}\;\Delta=\frac{\delta}{2\sigma_{e}}.$ (3.27)
From (3.26), it follows that if we reject $H_{0}$ when $|\mathcal{T}(n)|\geq u_{1-\alpha/2}$, then
$\lim_{n\to\infty}\textsf{P}(|\mathcal{T}(n)|\geq u_{1-\alpha/2}|H_{0})\leq\alpha.$
The test always controls the type I error rate but is conservative when $\sigma_{\mathcal{T}}<1$ ($\textsf{Pj}\big[(Y(1)-\mu_{1})/\rho+(Y(2)-\mu_{2})/(1-\rho)\big|\bm{\phi}(\bm{X})\big)\not\equiv 0$). Under the local alternatives, the asymptotic power is
$\lim_{n\to\infty}\textsf{P}(|\mathcal{T}(n)|\geq u_{\alpha}|H_{A}:\tau=\frac{\delta}{\sqrt{n}})=$ $\textsf{P}\left(\left|\frac{|\delta|}{2\sigma_{\tau}}+N(0,1)\right|\geq u_{1-\alpha/2}\frac{\sigma_{e}}{\sigma_{\tau}}\right)$
$\leq$ $\textsf{P}\left(\left|\frac{|\delta|}{2\sigma_{\tau}}+N(0,1)\right|\geq u_{1-\alpha/2}\right).$
If we replace $\widehat{\sigma}_{n,t}^{2}$ in (3.22) by the consistent estimators $\widehat{\widetilde{\sigma}}_{n,t}^{2}$ of $\widetilde{\sigma}_{t}^{2}$, $t=1,2$, and define the adjusted test statistic as
$\mathcal{T}_{adj}(n)=\frac{\overline{Y}_{n,1}-\overline{Y}_{n,2}}{\sqrt{\widehat{\widetilde{\sigma}}_{n,1}^{2}/N_{n,1}+\widehat{\widetilde{\sigma}}_{n,2}^{2}/N_{n,2}}}.$ (3.28)
then
$\lim_{n\to\infty}\textsf{P}(|\mathcal{T}_{adj}(n)|\geq u_{1-\alpha/2}|H_{0})=\alpha,$
$\lim_{n\to\infty}\textsf{P}(|\mathcal{T}_{adj}(n)|\geq u_{1-\alpha/2}|H_{A}:\tau=\frac{\delta}{\sqrt{n}})=\textsf{P}\left(\left|\frac{|\delta|}{2\sigma_{\tau}}+N(0,1)\right|\geq u_{1-\alpha/2}\right).$
The adjusted test restores the type I error rate and increases the power.
When $T_{i},\bm{\phi}(\bm{X}_{i})$, $i=1,\ldots,n$, are available in the analysis stage. We can obtain the consistent estimators as follow. Let $T_{i,1}=T_{i}$, $T_{i,2}=1-T_{i}$. We regress the residuals $T_{i,t}(Y_{i}-\overline{Y}_{n,1})$s over $T_{i,t}\bm{\phi}(\bm{X}_{i})$s by fitting the linear model
$T_{i,t}(Y_{i}-\overline{Y}_{n,t})=T_{i,t}\bm{\phi}(\bm{X}_{i})\bm{\alpha}_{t}^{\prime}+\zeta_{i,t},\;i=1,\ldots,n,$
and obtain the estimator $\widehat{\bm{\alpha}}_{n,t}$ of $\bm{\alpha}_{t}$ and the new residuals $\widehat{\zeta}_{i,t}=T_{i}(t)(Y_{i}-\overline{Y}_{n,t})-T_{i}(t)\bm{\phi}(\bm{X}_{i})\widehat{\bm{\alpha}}_{n,t}^{\prime}$, $i=1,\ldots,n$, $t=1,2$. Define
$\widehat{\widetilde{\sigma}}_{n,t}^{2}=\frac{1}{\sum_{i=1}^{n}T_{i}(t)-1}\sum_{i=1}^{n}\big(\widehat{\zeta}_{i,t}\big)^{2}+\frac{\rho(1-\rho)}{n-2}\sum_{i=1}^{n}\big(\bm{\phi}(\bm{X}_{i})[\widehat{\bm{\alpha}}_{n,1}-\widehat{\bm{\alpha}}_{n,2}]^{\prime}\big)^{2},$ (3.29)
$t=1,2$.
Theorem 3.2.
Under the conditions in Theorem 3.1,
$\widehat{\widetilde{\sigma}}_{n,t}^{2}\overset{P}{\to}\widetilde{\sigma}_{t}^{2},\;t=1,2.$
3.1 and 3.2 will follow from (2.7) and (2.8). When the stratified randomization is adopted, Theorems 3.1 and 3.2 remain true even when $\gamma=0$ by Theorem 2.4.
It is seen that the asymptotic variance $\vec{\sigma}_{Z}^{2}$ in (2.8) has the closed form $\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$ and does not exceed $\textsf{E}Z^{2}$ is important because in this case the classical test is valid and the variance is estimable so that the test can be adjusted to have the precise type I error rate.
When the stratified randomization is adopted, we always have $\vec{\sigma}_{Z}^{2}=\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$. This is the main reason why the classical test under the stratified randomization is always valid and can be adjusted in varies ways.
4 Proofs of the results
4.1 The properties of the CAR procedures
We write $\bm{\phi}_{n}=\bm{\phi}(\bm{X}_{n})$, $\bm{\phi}=\bm{\phi}(\bm{X})$. Let $\Sigma=\textsf{E}\left[\bm{\phi}^{\prime}\bm{\phi}\right]$. Without loss of generality, we assume $\Sigma$ is non-singular. Otherwise, there is an orthogonal matrix $U$ such that $\Sigma=U^{\prime}diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2},0,\ldots)U$, where $\sigma_{t}^{2}$s are positive numbers. Let $\widetilde{\bm{\phi}}=\bm{\phi}U^{\prime}$. Then $\textsf{E}[\widetilde{\bm{\phi}}^{\prime}\widetilde{\bm{\phi}}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2},0,\ldots)$, $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i}U$, $\|\bm{\Lambda}_{n}\|^{2}=\|\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i}\|^{2}$, $\langle\bm{\Lambda}_{n},\bm{\phi}_{n-1}\rangle=\langle\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i},\widetilde{\bm{\phi}}_{n+1}\rangle$. So, it is sufficient to consider $\widetilde{\bm{\phi}}=(\widetilde{\phi}^{(1)},\ldots,\widetilde{\phi}^{(q)})$. Since $\textsf{P}(\widetilde{\phi}^{(t)}=0,t=q_{1}+1,\ldots q)=1$, it is sufficient to consider the subverter $\widetilde{\bm{\psi}}=(\widetilde{\phi}^{(1)},\ldots,\widetilde{\phi}^{(q_{1})})$ with $\textsf{E}[\widetilde{\bm{\psi}}^{\prime}\widetilde{\bm{\psi}}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2})$ being non-singular.
Let $\mathscr{F}_{n}=\sigma(\bm{X}_{i},T_{i},Z_{i},W_{i},i=1,2,\ldots,n)$ be the history sigma generated by results of variables and allocations up to stage, $\ell_{n}=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\Phi}_{n}\rangle}{(n-1)^{\gamma}}\right)$. Then, $\textsf{P}(T_{n+1}=1|\mathscr{F}_{n},\bm{X}_{n+1})=\ell_{n+1}$ by (2.1). Let $r\in[2,r_{0}]$. Write $\bm{\psi}_{n}=(T_{i}-\rho)\bm{\phi}(\bm{X}_{n})$. Then $\bm{\Lambda}_{n+1}=\bm{\Lambda}_{n}+\bm{\psi}_{n}$. From the elementary inequality that
$\|\bm{u}+\bm{v}\|^{r}-\|\bm{u}\|^{r}\leq r\langle\bm{u},\bm{v}\rangle\|\bm{u}\|^{r-2}+c_{r}\left(\|\bm{v}\|^{r}+\|\bm{v}\|^{2}\|\bm{u}\|^{r-2}\right),\;r\geq 2,$ (4.30)
where $c_{r}$ is a constant which depends only on $r$, we have
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n},\bm{X}_{n+1}\right]-\|\bm{\Lambda}_{n}\|^{r}$
$\leq$ $\textsf{E}\left[r\langle\bm{\Lambda}_{n},\bm{\psi}_{n+1}\rangle\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\|\bm{\psi}_{n+1}\|^{r}+c_{r}\|\bm{\psi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\big|\mathscr{F}_{n},\bm{X}_{n+1}\right]$
$\leq$ $\textsf{E}\left[r(T_{n+1}-\rho)\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1})\rangle\|\bm{\Lambda}_{n}\|^{r-2}\big|\mathscr{F}_{n},\bm{X}_{n+1}\right]+c_{r}\left(\|\bm{\phi}_{n+1}\|^{r}+\|\bm{\phi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\right)$
$=$ $r(\ell_{n+1}-\rho)\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\left(\|\bm{\phi}_{n+1}\|^{r}+\|\bm{\phi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\right)$
$=$ $-r\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\cdot\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\left(\|\bm{\phi}_{n+1}\|^{r}+\|\bm{\phi}_{n+1}\|^{2}\|\bm{\Lambda}_{n}\|^{r-2}\right).$
Notice that there is a $\delta>0$, $|\ell(x)-\rho|\geq|x||l^{\prime}(0)|/2$ when $|x|\leq\delta$. While, when $|x|\geq\delta$, $|\ell(x)-\rho|\geq(\rho-\ell(\delta)\wedge(\ell(-\delta)-\rho)>0$. It follows that, there exists a contact $\lambda_{1}>0$ such that $|\ell(x)-\rho|\geq\lambda_{1}(|x|\wedge 1)$. Thus, $\textsf{E}\left[\left|\ell_{n+1}-\rho\right|\cdot\left|\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle\right|\big|\mathscr{F}_{n}\right]\geq\lambda_{1}n^{\gamma}\hbar(\bm{\lambda}_{n}/n^{\gamma})$, where $\hbar(\bm{x})=\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]$. Next, we show that here exists a constant $\lambda_{2}>0$ such that $\hbar(\bm{x})\geq\lambda_{2}(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)$. Under the assumption that $\Sigma=\textsf{E}[\bm{\phi}^{\prime}\bm{\phi}]$ is non-singular, we have $\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]\neq 0$ for all $\bm{x}\neq\bm{0}$. In fact, if $\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]=0$, then $\textsf{P}(\langle\bm{x},\bm{\phi}\rangle=0)=1$ which implies that $\bm{x}\Sigma\bm{x}^{\prime}=\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\right]=0$. Thus, we must have $\bm{x}=\bm{0}$ since $\Sigma$ is non-singular.
Notice that $\hbar(\bm{x})$ is a continuous function. Let $\lambda_{0}=\inf\{\textsf{E}\left[\left|\langle\bm{x},\bm{\phi}\rangle\right|\right]:\|\bm{x}\|=1\}$. Then $\lambda_{0}>0$. Since $\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|\leq(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)(\|\bm{\phi}\|^{2}+\|\bm{\phi}\|)$, by the dominated convergence theorem, as $\|\bm{x}\|\to 0$ with $\bm{x}/\|\bm{x}\|\to\bm{x}_{0}$,
$\lim_{\|\bm{x}\|\to 0}\frac{\hbar(\bm{x})}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}=$ $\textsf{E}\left[\lim_{\|\bm{x}\|\to 0}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}\right]=\textsf{E}\left[\lim_{\|\bm{x}\|\to 0}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}}{\|\bm{x}\|^{2}}\right]$
$=$ $\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|^{2}\right]\geq\big(\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|\right]\big)^{2}=\lambda_{0}^{2}>0;$
and, as $\|\bm{x}\|\to\infty$ with $\bm{x}/\|\bm{x}\|\to\bm{x}_{0}$,
$\lim_{\|\bm{x}\|\to\infty}\frac{\hbar(\bm{x})}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}=$ $\textsf{E}\left[\lim_{\|\bm{x}\|\to\infty}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|^{2}\wedge\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|^{2}\wedge\|\bm{x}\|}\right]=\textsf{E}\left[\lim_{\|\bm{x}\|\to\infty}\frac{\left|\langle\bm{x},\bm{\phi}\rangle\right|}{\|\bm{x}\|}\right]$
$=$ $\textsf{E}\left[\left|\langle\bm{x}_{0},\bm{\phi}\rangle\right|\right]\geq\lambda_{0}>0.$
Thus, there exists $\lambda_{2}>0$ such that $\hbar(\bm{x})\geq\lambda_{2}(\|\bm{x}\|^{2}\wedge\|\bm{x}\|)$. It follows that
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}$
$\leq$ $-r\lambda_{1}n^{\gamma}\hbar\left(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\right)\cdot\|\bm{\Lambda}_{n}\|^{r-2}+c_{r}\left(\beta_{r}+\beta_{2}\|\bm{\Lambda}_{n}\|^{r-2}\right),$
$\leq$ $-r\lambda_{1}\lambda_{2}\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)+c_{r}\left(\beta_{r}+\beta_{2}\|\bm{\Lambda}_{n}\|^{r-2}\right),$ (4.31)
where $\beta_{r}=\textsf{E}[\|\bm{\phi}(\bm{X})\|^{r}]$. Hence,
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\right]-\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]$
$\leq$ $-r\lambda_{1}\lambda_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)\right]+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-2}\right]\right)$ (4.32)
$\leq$ $c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-2}\right]\right).$ (4.33)
Taking $r=2$ in (4.33) yields $\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]\leq 2c_{2}\beta_{2}n$. Thus, $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq(\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}])^{r/2}\leq C_{r}n^{r/2}$ for all $n$ and $r\in(0,2]$. By (4.33) and the induction,
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq C_{r}n^{r/2},\;\;n\geq 1,r\in(0,r_{0}].$ (4.34)
Let $E=\{\|\bm{\Lambda}_{n}\|\leq n^{\gamma}\}$ and $Q_{n,r}=\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)\right]$. Then
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}I_{E}\right]\leq n^{\gamma}Q_{n,r}\;\text{ and }\;\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}I_{E^{c}}\right]\leq Q_{n,r}.$
For $\frac{1}{p}+\frac{1}{q}=1$, $p,q>1$, by the Hölder inequality we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}I_{E^{c}}\right]=\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{\frac{r-1}{p}}\|\bm{\Lambda}_{n}\|^{\frac{r-1}{q}+1}I_{E^{c}}\right]$
$\leq$ $\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1}I_{E^{c}}\right]\right)^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}I_{E^{c}}\right]\right)^{1/q}\leq Q_{n,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}\right]\right)^{1/q}.$
It follows that
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq n^{\gamma}Q_{n,r}+Q_{n,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r-1+q}\right]\right)^{1/q}.$
Now, let $m:=m_{n}\in\{0,1,\ldots,n\}$ be the last one for which
$r\lambda_{1}\lambda_{2}Q_{m,r}+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)\geq 0.$
Then
$Q_{m,r}\leq C_{r}\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
and, by (4.32) and (4.33),
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
$\leq$ $m^{\gamma}Q_{m,r}+Q_{m,r}^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+q}\right]\right)^{1/q}+c_{r}\left(\beta_{r}+\beta_{2}\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)$
$\leq$ $C_{r}\left(m^{\gamma}\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)+\left(1+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-2}\right]\right)^{1/p}\left(\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1+q}]\right)^{1/q}\right).$ (4.35)
Proof of Theorem 2.1. Let $r=2$ and $r-1+q=r_{0}$. Notice $\textsf{E}[\|\bm{\Lambda}_{m}\|^{r_{0}}]\leq C_{r_{0}}m^{r_{0}/2}$. We obtain
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]\leq C\left(n^{\gamma}+n^{1+\frac{2-r_{0}}{2(r_{0}-1)}}\right)=o(n),$
by (4.1). Thus, (2.5) is proven.
Notice that
$\textsf{E}|(\ell_{n}-1)Z_{n}|\leq$ $cM\textsf{E}\left[\frac{\|\bm{\Lambda}_{n-1}\|\cdot\|\bm{\phi}_{n}\|}{(n-1)^{\gamma}}\right]+\textsf{E}[|Z|I\{|Z|>M\}]$
$=$ $cM\frac{\textsf{E}\left[\|\bm{\Lambda}_{n-1}\|\right]\textsf{E}[\|\bm{\phi}\|]}{(n-1)^{\gamma}}+\textsf{E}[|Z|I\{|Z|>M\}]$
$\to$ $0\text{ as }n\to\infty\text{ and then }M\to\infty,$
since
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]=O\left(n^{\gamma/2}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}\right)=o(n^{\gamma}),\;\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}.$
(2.6) is proven.
For (2.7), let $\Delta M_{i,Z}=:(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]$, $M_{n,Z}=\sum_{i=1}^{n}\Delta M_{i,Z}$. Notice that $\{\Delta M_{i,Z},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences. Since $|\Delta M_{i,Z}|\leq|Z_{i}|+\textsf{E}|Z|$, $\{\Delta M_{i,Z},i=1,2,\ldots\}$ is uniformly integrable. Thus, $\frac{1}{n}\textsf{E}[|M_{n,Z}|\to 0$ (Hall and Heyde, , 1980, Theorem 2.2). On the other hand,
$\textsf{E}\Big|\sum_{i=2}^{n}\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]\Big|\leq\sum_{i=2}^{n}\textsf{E}[|(\ell_{i}-\rho)Z_{i}|]=o(n)$
by (2.6). Hence, (2.7) is proven.
For (2.8), we denote $\Delta M_{i,\widetilde{Z}}=:(T_{i}-\rho)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]$. Suppose $r_{0}\geq 4$ and $\textsf{E}[Z^{2}]<\infty$. Then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}+\langle\bm{\Lambda}_{n},\bm{x}_{0}\rangle$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+\langle\bm{\Lambda}_{n},\bm{x}_{0}\rangle+\sum_{i=2}^{n-1}\textsf{E}[(\ell_{i+1}-\rho)\widetilde{Z}_{i+1}|\bm{\Lambda}_{i}].$ (4.36)
We notice that
$|\ell(x)-\rho-\ell^{\prime}(0)x|\leq cx^{2}.$
[…]
$\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]\right|$
$=$ $\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]-\ell^{\prime}(0)h_{\widetilde{Z}}\Big(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\Big)\right|\leq C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{2}\|\bm{\Phi}\|^{2}|\widetilde{Z}|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]$
$\leq$ $C\frac{\|\bm{\Lambda}_{n}\|^{2}}{n^{2\gamma}}\Big(\textsf{E}\widetilde{Z}^{2}\Big)^{1/2}\Big(\textsf{E}\|\bm{\phi}\|^{4}\Big)^{1/2}\leq C_{0}\big(\textsf{E}Z^{2}\big)^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{2}}{n^{2\gamma}}.$ (4.37)
Also, since $\bm{x}_{0}\Sigma\bm{x}_{0}=\textsf{E}(\textsf{Pj}[Z|\bm{\phi}])^{2}\leq\textsf{E}Z^{2}$, we have $\|\bm{x}_{0}\|\leq C_{0}(\textsf{E}Z^{2})^{1/2}$. Hence, we have
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+\theta_{n}\left(\|\bm{\Lambda}_{n}\|+\sum_{i=2}^{n-1}\frac{\|\bm{\Lambda}_{i}\|^{2}}{i^{2\gamma}}\right),\;\;|\theta_{n}|\leq C_{0},$ (4.38)
where $C_{0}$ does not depend on $Z$.
From (2.5), it follows that
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O\Big(n^{\frac{\gamma}{2}}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}+n^{1-\gamma}+n^{\frac{3}{2}+\frac{1}{2(r_{0}-1)}-2\gamma}\Big)$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+o(n^{1/2})\text{ in }L_{1}$
by the facts that $\frac{\gamma}{2}<\frac{1}{2}$, $\frac{1}{4}+\frac{1}{4(r_{0}-1)}<\frac{1}{2}$, $1-\gamma<\frac{1}{2}$ and $\frac{3}{2}+\frac{1}{2(r_{0}-1)}-2\gamma<\frac{1}{2}$.
Notice that $\{(\Delta M_{i,\tilde{Z}},W_{i}),\mathscr{F}_{i},i=1,2,\ldots\}$ is sequence of martingale differences. We have $|\Delta M_{i,\tilde{Z}}|\leq|\widetilde{Z}_{i}|+\textsf{E}|\widetilde{Z}_{i}|$,
$\frac{1}{n}\sum_{i=1}^{n}\textsf{E}\left[(\Delta M_{i,\tilde{Z}})^{2}I\{\big|(\Delta M_{i,\tilde{Z}})^{2}\geq\epsilon n\}\right]$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\textsf{E}\left[(|\widetilde{Z}_{i}|+\textsf{E}|\widetilde{Z}|)^{2}I\{\big|(|\widetilde{Z}_{i}|+\textsf{E}|\widetilde{Z}|)^{2}\geq\epsilon n\}\right]$
$=$ $\textsf{E}\left[(|\widetilde{Z}|+\textsf{E}|\widetilde{Z}|)^{2}I\{\big|(|\widetilde{Z}|+\textsf{E}|\widetilde{Z}|)^{2}\geq\epsilon n\}\right]\to 0,$
$\frac{1}{n}\sum_{i=1}^{n}\textsf{E}\left[W_{i}^{2}I\{W_{i}^{2}\geq\epsilon n\}\right]=\textsf{E}\left[W^{2}I\{W^{2}\geq\epsilon n\}\right]\to 0,$
i.e., the Linderberg condition is satisfied. Further, and $\textsf{E}[W_{i}^{2}|\mathscr{F}_{i-1}]=\textsf{E}W^{2}$,
$\textsf{E}\left[(\Delta M_{i,\tilde{Z}})^{2}\big|\mathscr{F}_{i-1}\right]=$ $\textsf{E}[((1-2\rho)\ell_{i}+\rho^{2})\tilde{Z}_{i}^{2}|\bm{\Lambda}_{i-1}]-\big(\textsf{E}[(\ell_{i}-\rho)\tilde{Z}_{i}|\bm{\Lambda}_{i-1}]\big)^{2}$
$\to$ $(1-\rho)\rho\textsf{E}[\tilde{Z}_{i}^{2}]\text{ in }L_{1},$
$\textsf{E}\left[\Delta M_{i,\tilde{Z}}W_{i}\big|\mathscr{F}_{i-1}\right]=$ $\textsf{E}[(\ell_{i}-\rho)\tilde{Z}_{i}W_{i}|\bm{\Lambda}_{i-1}]\to 0\text{ in }L_{1},$
by applying (2.6) to $\tilde{Z}^{2}_{n}$ and $\tilde{Z}_{n}W_{n}$.
It follows that
$\frac{\textsf{E}\big(M_{n,\widetilde{Z}},\sum_{i=1}^{n}W_{i}\big)^{\otimes 2}}{n}\to diag(\vec{\sigma}^{2},\textsf{Var}(W))\text{ and }$ (4.39)
$\left(\frac{M_{n,\widetilde{Z}}}{\sqrt{n}},\frac{\sum_{i=1}^{n}W_{i}}{\sqrt{n}}\right)\overset{d}{\to}N(0,\sigma^{2})\times N(0,\textsf{Var}(W)),$
by the central limit theorem for martingales(Hall and Heyde, , 1980, Corollary 3.1). Hence, (2.8) holds. $\Box$
Proof of Theorem 2.2. We will use the induction method to obtain the moment estimates of $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]$ by inducting both in $n$ and $r$. We assume
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{s}]\leq C_{s}n^{\beta s/2}\text{ for all integers }n,s\geq 0,$ (4.40)
and
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r-2}]\leq C_{r}n^{\alpha(r-2)/2}\text{ for all }n,$ (4.41)
where $0<\beta/2<\alpha\leq\beta\leq 2$.
Let $q$ be an integer large enough such that
$\frac{(\beta-\alpha)r-\beta}{2q}+\frac{\beta}{2}-\alpha<0.$
Then by (4.1),
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq$ $C_{r}\left(m^{\frac{\alpha(r-2)}{2}+\gamma}+m^{\frac{\alpha(r-2)}{2p}+\frac{\beta(r-1+q)}{2q}}\right)$
$\leq$ $C_{r}\left(n^{\frac{\alpha r}{2}+\gamma-\alpha}+n^{\frac{\alpha r}{2}+\frac{(\beta-\alpha)r-\beta}{2q}+\frac{\beta}{2}-\alpha}\right)$
$\leq$ $C_{r}n^{\alpha r/2}\;\;\text{ for all }n,$ (4.42)
whenever $\alpha\geq\gamma$.
Obviously, (4.40) holds for $\beta=1$ by (4.34), and (4.41) is true for $r=2$. So, by (4.1) and the induction, (4.41) holds for all integer $r\geq 2$ whenever $\alpha\geq\gamma$ and $\alpha>1/2$, which implies that (4.40) holds for any $\beta$ with $\beta\geq\gamma$ and $\beta>1/2$. Repeating the above procedure $i$ times concludes that (4.40) holds for any $\beta$ with $\beta\geq\gamma$ and $\beta>1/2^{i-1}$. Stooping the procedure when $1/2^{i-1}<\gamma$, we conclude that
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq C_{r}n^{\gamma r/2}\;\text{ for all integers }n\text{ and }r\geq 2.$
(2.9) is proven.
When $\textsf{E}|Z|<\infty$, the proofs of (2.6) and (2.7) are the same as those in Theorem 2.1 since $\textsf{E}\|\bm{\Lambda}_{n}\|=o(n^{\gamma})$.
Suppose $\textsf{E}[Z^{2}]<\infty$. Then ((ii)) follows from (2.9) and (4.38), and, (2.8) follows from ((ii)) and (4.38). Notice
$\textsf{E}M_{n,\widetilde{Z}}^{2}=\sum_{i=1}^{n}\textsf{E}(\Delta M_{i,\tilde{Z}})^{2}\leq n\textsf{E}\widetilde{Z}^{2}\leq n\textsf{E}Z^{2}.$
By (4.38) and (2.9), (2.12) holds. $\Box$
Proof of Remark 2.2. Notice
$\left|\ell(x)-\rho-\ell^{\prime}(0)x-\frac{1}{2}\ell^{\prime\prime}(0)x^{2}\right|\leq c|x|^{3}.$
Similar to (4.1), we have
$\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]-\ell^{\prime}(0)h_{\widetilde{Z}}\Big(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\Big)-\frac{1}{2}\ell^{\prime\prime}(0)\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\textsf{E}\left[\bm{\phi}^{\prime}\bm{\phi}\widetilde{Z}\right]\frac{\bm{\Lambda}_{n}^{\prime}}{n^{\gamma}}\right|$
$\leq$ $C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{3}\|\bm{\Phi}\|^{3}|\widetilde{Z}|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq C_{0}(\textsf{E}Z^{2})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{3}}{n^{3\gamma}}.$
Notice that $h_{\widetilde{Z}}\big(\bm{x}\big)\equiv\bm{0}$ and $\ell^{\prime\prime}(0)\textsf{E}[\bm{\phi}^{\prime}\bm{\phi}\widetilde{Z}]=\bm{0}$. It follows that
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+\theta_{n}(\textsf{E}Z^{2})^{1/2}\left(\|\bm{\Lambda}_{n}\|+\sum_{i=2}^{n-1}\frac{\|\bm{\Lambda}_{i}\|^{3}}{i^{3\gamma}}\right)\;(|\theta_{n}|\leq C_{0})$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O(n^{\gamma/2}+n^{1-3\gamma/2})\text{ in }L_{r},$
where $C_{0}$ does not depend on $Z$.
Proof of Theorem 2.3. Notice $\Phi(u_{\rho/2})<\rho<\Phi(u_{(\rho+1)/2})$. Let $K>2$. By the continuity, there exists $\delta>0$ such that
$\ell(x)=\rho-\lambda x,\;\;|x|\leq\delta.$
When $|x|\geq\delta$, we have $|\ell(x)-\rho|\leq 2\leq 2\delta^{-K}|x|^{K}$, $|x|\leq\delta^{-K+1}|x|^{K}$. So, there exists a constant $C_{K}$ such that
$|\ell(x)-\rho+\lambda x|\leq c_{K}|x|^{K}.$
Similar to (4.1), we have
$\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]\right|=\left|\textsf{E}\left[(\ell_{n+1}-\rho)\widetilde{Z}_{n+1}\big|\bm{\Lambda}_{n}\right]+\lambda h_{\widetilde{Z}}\Big(\frac{\bm{\Lambda}_{n}}{n^{\gamma}}\Big)\right|$
$\;\;\leq c_{K}\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{K}\|\bm{\Phi}\|^{K}|Z|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq C_{K}(\textsf{E}Z^{2})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{K}}{n^{K\gamma}}.$
Choose $K>2$ such that $K\gamma/2>1$. By (2.9),
$\sum_{i=1}^{n-1}\frac{\|\bm{\Lambda}_{i}\|^{K}}{i^{K\gamma}}\leq C\sum_{i=1}^{\infty}i^{-K\gamma/2}<\infty\;\;\text{ in }L_{r}.$
The proof is completed by (4.1) and (2.9). $\Box$
Proof of Theorem 2.4. For (i), notice that (4.1) still holds for $\gamma=0$. When $\|\bm{\Lambda}_{n}\|\geq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$,
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+c_{r}\beta_{r}.$
When $\|\bm{\Lambda}_{n}\|\leq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$, $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+C_{r}$. Thus, exist $a_{r},b_{r}>0$ such that $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-a_{r}\|\bm{\Lambda}_{n}\|^{r-1}+b_{r}.$ It follows that
$\textsf{E}[\|\bm{\Lambda}_{n+1}\|^{r}]-\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq$ $-a_{r}\textsf{E}[\|\bm{\Lambda}_{n}\|^{r-1}]+b_{r}$ (4.43)
$\leq$ $b_{r}.$ (4.44)
From (4.44), it follows that $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r_{0}}]\leq nb_{r_{0}}$.
Now, let $m:=m_{n}\in\{0,1,\ldots,n\}$ be the last one for which $a_{r}\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1}]+b_{r}\geq 0.$ Let $r-1+q=r_{0}$, $1/p+1/q=1$. Then
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq$ $\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+b_{r}=b_{r}+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{\frac{r-1}{p}}\|\bm{\Lambda}_{m}\|^{\frac{r-1}{q}+1}\right]$
$\leq$ $b_{r}+\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1}\right]\right)^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+q}\right]\right)^{1/q}$
$\leq$ $b_{r}+(b_{r}/a_{r})^{1/p}\left(b_{r_{0}}m\right)^{1/q}\leq C_{r,r_{0}}n^{1/q}.$ (4.45)
Letting $r=2$ and $q=r_{0}-1$ yields $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{\frac{1}{r_{0}-1}})$. When $r_{0}\geq 3$. Let $r=(1+r_{0})/2$. Then $r\geq 2$, $q=r$. Thus, $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]=O(n^{1/r})$, and so $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{2/r^{2}})=O(n^{\frac{8}{(r_{0}+1)^{2}}}$. The proof of (i) is completed.
For (ii), suppose $\textsf{E}[Z|\bm{\phi}]=\langle\bm{\phi},\bm{z}_{0}\rangle\in Span\{\bm{\phi}\}$. Let $\widetilde{Z}=Z-\textsf{E}[Z|\bm{\phi}]$. Then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}+\langle\bm{\Lambda}_{n},\bm{z}_{0}\rangle,$
and $\{(T_{i}-\rho)\widetilde{Z}_{i},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences with $|(T_{i}-\rho)\widetilde{Z}_{i}|\leq|Z_{i}|+\|\bm{z}_{0}\|\cdot\|\bm{\phi}_{i}\|$. Thus, the martingale differences are uniformly integrable which implies
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}}{n}\to 0\text{ in }L_{1}$
(Hall and Heyde, , 1980, Theorem 2.22). (2.7) is proved by (i).
Suppose $\textsf{E}Z^{2}<\infty$. In this case $\langle\bm{\phi},\bm{z}_{0}\rangle=\textsf{Pj}[Z|\bm{\phi}]$ and thus $|\langle\bm{\Lambda}_{n},\bm{z}_{0}\rangle|\leq C_{0}(\textsf{E}Z^{2})^{1/2}\|\bm{\Lambda}_{n}\|$ as we have shown. For the martingale $\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}$, we have
$\textsf{E}\big(\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}\big)^{2}\leq n\textsf{E}\widetilde{Z}^{2}\leq n\textsf{E}Z^{2}.$
(ii) is proven.
For (iii), by the central limit theorem(Hall and Heyde, , 1980, Theorem 3.2), it is sufficient to show that
$\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\rho)^{2}\widetilde{Z}_{i}^{2}\to\rho(1-\rho)\textsf{E}\widetilde{Z}^{2}\;\;\text{ in }L_{1}.$ (4.46)
Notice
$(T_{i}-\rho)^{2}\widetilde{Z}_{i}^{2}=\rho(1-\rho)\textsf{E}\widetilde{Z}^{2}+\rho(1-\rho)(\widetilde{Z}_{i}^{2}-\textsf{E}\widetilde{Z}^{2})+(1-2\rho)(T_{i}-\rho)\widetilde{Z}_{i}^{2}.$
It is obvious that
$\frac{\sum_{i=1}^{n}(\widetilde{Z}_{i}^{2}-\textsf{E}\widetilde{Z}^{2})}{n}\to 0\text{ in }L_{1}.$
When $\rho\neq 1/2$, we have
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}^{2}}{n}\to 0\text{ in }L_{1},$
by (ii) and noticing the assumption that $\textsf{E}[\widetilde{Z}^{2}|\bm{\phi}]\in Span\{\bm{\phi}\}$ . Hence, (4.46) holds.
When $(Z-\textsf{E}[Z|\bm{\phi}])W$ is also in $Span\{\bm{\phi}\}$. Then
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}W_{i}}{n}\to 0\text{ in }L_{1}$
by (ii) again. Thus,
$\sum_{i=1}^{n}\big((T_{i}-\rho)\widetilde{Z}_{i},W_{i}\big)^{\otimes 2}\to diag\big(\vec{\sigma}_{Z}^{2},\textsf{E}W^{2}\big)\text{ in }L_{1}.$
Applying the central limit theorem to martingale differences $\Big\{\big((T_{i}-\rho)\widetilde{Z}_{i},W_{i}\big),\mathscr{F}_{i},i=1,2,\ldots\Big\}$ yields (2.8). The proof is completed. $\Box$
4.2 The properties of the test
Proof of Theorem 3.1. (3.23) is trivial by applying (2.7). For (3.24), we let $e_{t}=Y(t)-\mu_{t}$, $\tilde{e}_{t}=e_{t}-\textsf{Pj}[e_{t}|\bm{\phi}]$, $e_{i,t}=Y_{i}(t)-\mu_{t}$, $D_{n,t}=\sum_{i=1}^{n}(T_{i}-\rho)e_{i,t}$, $t=1,2$, $D_{n}=\sum_{i=1}^{n}(T_{i}-\rho)$. Then $D_{n}=O_{p}(n^{1/2})$, $D_{n,t}=O_{p}(n^{1/2})$, $\sum_{i=1}^{n}e_{i,t}=O_{P}(n^{1/2})$,$t=1,2$, by (2.8). Thus,
$\widehat{\tau}_{n}=:$ $\overline{Y}_{n,1}-\overline{Y}_{n,1}=\tau+\frac{\sum_{i=1}^{n}e_{i,1}+\frac{D{{}_{n,1}}}{\rho}}{n\big(1+\frac{D_{n}}{n\rho}\big)}-\frac{\sum_{i=1}^{n}e_{i,2}-\frac{D{{}_{n,2}}}{1-\rho}}{n\big(1-\frac{D_{n}}{n(1-\rho)}\big)}$
$=$ $\tau+\frac{\frac{D{{}_{n,1}}}{\rho}+\frac{D{{}_{n,2}}}{1-\rho}+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{1/2})$
$=$ $\tau+\frac{\sum_{i=1}^{n}(T_{i}-\rho)\big(\frac{e{{}_{i,1}}}{\rho}+\frac{e{{}_{i,2}}}{1-\rho}\big)+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{1/2}).$
By (2.8), it follows that
$\sqrt{n}(\widehat{\tau}_{n}-\tau)\overset{d}{\to}N(0,\sigma_{\tau}^{2})$
with
$\sigma_{\tau}^{2}=$ $\rho(1-\rho)\textsf{E}\Big(\frac{\tilde{e}_{1}}{\rho}+\frac{\tilde{e}_{2}}{1-\rho}\Big)^{2}+\textsf{E}(e_{1}-e_{2})^{2}$
$=$ $\rho(1-\rho)\textsf{E}\Big(\frac{e_{1}}{\rho}+\frac{e_{2}}{1-\rho}\Big)^{2}+\textsf{E}(e_{1}-e_{2})^{2}-\rho(1-\rho)\textsf{E}\Big(\textsf{Pj}[\frac{e_{1}}{\rho}+\frac{e_{2}}{1-\rho}|\bm{\phi}]\Big)^{2}$
$=$ $\frac{\textsf{E}e_{1}^{2}}{\rho}+\frac{\textsf{E}e_{2}^{2}}{1-\rho}-\rho(1-\rho)\textsf{E}\Big(\textsf{Pj}[\frac{e_{1}}{\rho}+\frac{e_{2}}{1-\rho}|\bm{\phi}]\Big)^{2}.$
Also,
$\sigma_{\tau}^{2}=$ $\rho(1-\rho)\textsf{E}\Big(\frac{\tilde{e}_{1}}{\rho}+\frac{\tilde{e}_{2}}{1-\rho}\Big)^{2}+\textsf{E}(\tilde{e}_{1}-\tilde{e}_{2})^{2}+\textsf{E}\big(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}]\big)^{2}$
$=$ $\frac{\textsf{E}\tilde{e}_{1}^{2}+\rho(1-\rho)\textsf{E}(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}])^{2}}{\rho}+\frac{\textsf{E}\tilde{e}_{2}^{2}+\rho(1-\rho)\textsf{E}(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}])^{2}}{1-\rho}.$
(3.24) is proven. For (3.26), it is sufficient to notice that
$\frac{n\widehat{\sigma}_{n,1}^{2}}{N_{n,1}}+\frac{n\widehat{\sigma}_{n,2}^{2}}{N_{n,2}}=\frac{\widehat{\sigma}_{n,1}^{2}}{1+\frac{D_{n}}{n\rho}}+\frac{\widehat{\sigma}_{n,2}^{2}}{1-\frac{D_{n}}{n(1-\rho)}}\overset{P}{\to}\frac{\sigma_{1}^{2}}{\rho}+\frac{\sigma_{2}^{2}}{1-\rho}.\;\;\Box$
Proof of Theorem 3.2. For simplifying the proof, without loss of generality we assume that $\Sigma=\textsf{E}[\bm{\phi}^{\otimes 2}]$ is non-singular. Let $\bm{\alpha}_{t}=\textsf{E}[\bm{\phi}e_{t}]\Sigma^{-1}$. Then $\bm{\phi}(\bm{X})\bm{\alpha}_{t}^{\prime}=\textsf{Pj}[e_{t}|\bm{\phi}]$. By noting $\overline{Y}_{n,t}-\mu_{t}\overset{P}{\to}0$, we have $\widehat{\zeta}_{i}=Y_{i}-\underline{\bm{X}}_{i}\widehat{\theta}^{\prime}-\bm{\phi}(\bm{X}_{i})\widehat{\bm{\alpha}}^{\prime}$ with
$\widehat{\bm{\alpha}}_{n,1}=\sum_{i=1}^{n}T_{i}\bm{\phi}_{i}(Y_{i}(1)-\overline{Y}_{n,1})\left[\sum_{i=1}^{n}T_{i}\bm{\phi}_{i}^{\otimes 2}\right]^{-1}$
$=$ $\frac{1}{n\rho}\sum_{i=1}^{n}T_{i}\bm{\phi}_{i}(Y_{i}(1)-\mu_{1})\Sigma^{-1}+o_{P}(1)\overset{P}{\to}\textsf{E}\left[\bm{\phi}e_{1}\right]\Sigma^{-1}=\bm{\alpha}_{1},$
by (2.7). Thus
$\frac{1}{\sum_{i=1}^{n}T_{i}(1)-1}\sum_{i=1}^{n}\big(\widehat{\zeta}_{i,1}\big)^{2}=\frac{1}{N_{n,1}-1}\sum_{i=1}^{n}T_{i}\big(Y_{i}(1)-\overline{Y}_{n,1}-\bm{\phi}_{i}\widehat{\bm{\alpha}}_{n,1}^{\prime}\big)^{2}$
$=$ $\frac{1}{n\rho}\sum_{i=1}^{n}T_{i}\big(e_{i,1}-\bm{\phi}_{i}\bm{\alpha}_{1}^{\prime}\big)^{2}+o_{P}(1)$
$\overset{P}{\to}\textsf{E}\big(e_{1}-\bm{\phi}\bm{\alpha}_{1}^{\prime}\big)^{2}=\textsf{E}\Big(Y(1)-\mu_{1}-\textsf{Pj}\big[Y(1)-\mu_{1}\big|\bm{\phi}\big]\Big)^{2},$
by (2.7) again.
Similarly, $\widehat{\bm{\alpha}}_{n,2}\overset{P}{\to}\bm{\alpha}_{2}$,
$\frac{1}{\sum_{i=1}^{n}T_{i}(2)-1}\sum_{i=1}^{n}\big(\widehat{\zeta}_{i,2}\big)^{2}\overset{P}{\to}\textsf{E}\Big(Y(2)-\mu_{2}-\textsf{Pj}\big[Y(2)-\mu_{2}\big|\bm{\phi}\big]\Big)^{2},$
and
$\frac{1}{n-2}\sum_{i=1}^{n}\big(\bm{\phi}_{i}[\widehat{\bm{\alpha}}_{n,1}-\widehat{\bm{\alpha}}_{n,2}]^{\prime}\big)^{2}=\frac{1}{n}\sum_{i=1}^{n}\big(\bm{\phi}_{i}[\bm{\alpha}_{1}-\bm{\alpha}_{2}]^{\prime}\big)^{2}+o_{P}(1)$
$\overset{P}{\to}\textsf{E}\big(\bm{\phi}[\bm{\alpha}_{1}-\bm{\alpha}_{2}]^{\prime}\big)^{2}=\textsf{E}\Big(\textsf{Pj}\big[Y(1)-\mu_{1}-(Y(2)-\mu_{2})\big|\bm{\phi}(\bm{X})\big]\Big)^{2}.$
The proof is completed. $\Box$